OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point:(2,-6) ; Slope:-3/4

OpenStudy (anonymous):

you can use the formula, \[(y-y1) = m (x-x1)\] wherein m is your slope and 2 is x1 and -6 is y1.

OpenStudy (anonymous):

is it y-6=-3/4(x+2)?

OpenStudy (pawanyadav):

Y+6=-3/4(x-2)

OpenStudy (anonymous):

it should be y+6 and x-2 be careful with the signs and i think you have to expand it more :)

OpenStudy (anonymous):

thanks